The table shows the costs of two milk producers.

 

Cost per litre

Firm X $9

Firm Y $7

 

The price received by producers is $10 per litre. Both firms have been given quotas allowing them to produce 200 litres per day. Firm X sells its quota to firm Y.

Assuming constant costs of production and zero costs of entry and exit, calculate the price range which firm Y had to pay (per day) to buy X’s quota.


1
Expert's answer
2021-12-07T21:20:27-0500

"Solution"

"Firm \ x=9 \times\ 200 = 1800\\
 \ y=7 \times\ 200=1400\\
Total \ costs=3200\\
 Price \ given = 10\\
Each \ 200\times\ 2 \times\ 10=4000\\
Therefore=4000-3200=\$ 800"